OpenStudy (anonymous):

\[\int\frac{1}{\sqrt{sinx}} dx\]

OpenStudy (lgbasallote):

zomg

OpenStudy (anonymous):

lol, whats wrong

OpenStudy (lgbasallote):

algebraic sub....

OpenStudy (anonymous):

try it then :P

Parth (parthkohli):

\[u = \sqrt{\sin x}\]\[u^2 = \sin x\]\[x = \sin^{-1}(u^2)\]Is that how you start?

OpenStudy (lgbasallote):

this looks like one of @Mimi_x3's old integrals

OpenStudy (anonymous):

you will get into a dead end

Parth (parthkohli):

Right, what do I have to do?

OpenStudy (anonymous):

its dirty integral :P

Parth (parthkohli):

I dislike the taste of ugly integrals :)

OpenStudy (anonymous):

hehe uglier than ugly integrals :P

OpenStudy (anonymous):

substitute sin x= t^2

OpenStudy (anonymous):

it is an integral that does not have a closed form..

OpenStudy (lgbasallote):

reduction formula

OpenStudy (lgbasallote):

wallis formula

OpenStudy (anonymous):

lol, do you know what it is igba

OpenStudy (lgbasallote):

riemann sum

OpenStudy (lgbasallote):

no idea. fancy words?

OpenStudy (anonymous):

or use sqrt(cosec x) = 1/2[{ (sqrt(cosec x) + cot (x))} + { sqrt(cosec x) - sqrt cot (x)}]

OpenStudy (anonymous):

last time i checked Riemann sum was used for definite integral :P

OpenStudy (anonymous):

sqrt(cosec x) = 1/2[ {sqrt(cosec x) + sqrtcot (x))} + { sqrt(cosec x) - sqrt cot (x)}]

Parth (parthkohli):

How about just saying the following? :P\[\ln|\sqrt{\sin x}|\]

OpenStudy (anonymous):

lol, that is not integrating properly.. i iwsh it was that easy :p

OpenStudy (anonymous):

sqrt(cosec x) = 1/2[{ (sqrt(cosec x) + cot (x))} + { sqrt(cosec x) - sqrt cot (x)}].. use this and integrate the rhs ...i.e sum of two integrands so integrate them separately

OpenStudy (anonymous):

but its an integral that does not have a closed form; can you do it like that?

OpenStudy (anonymous):

There's actual theory dedicated to this... it's not definite: http://en.wikipedia.org/wiki/Elliptic_integral

OpenStudy (shubhamsrg):

sinx = t^2 =>cos x = sqrt(1-t^4) now dx = 2t dt / (sqrt(1-t^4)) making that substituion, we have 2 integral ( dt / sqrt(1-t^4) ) does this help ?

OpenStudy (anonymous):

(definite as in 'closed form', exact expression)

OpenStudy (anonymous):

well does it mean it cant be solved analytically?